OpenStudy (anonymous):

find the volume of the tetrahedron bounded by the planes x+2y+z=2 , x=2y , x=0 , z=0

OpenStudy (anonymous):

\[\int\limits_{x=0}^{2} \int\limits_{y=0}^{x/2} \int\limits_{z=0}^{2-x-2y}dzdydx\] Try to draw a sketch of the solid.

OpenStudy (anonymous):

ok. i got it. thanks